OpenStudy (anonymous):

67/100 in lowest terms as a fraction

OpenStudy (anonymous):

67/100

OpenStudy (anonymous):

.6

OpenStudy (anonymous):

\[\Xi \zeta \epsilon \delta \gamma \xi \Phi \Psi \Omega \left[\begin{matrix}\left(\begin{matrix}x \\ z\end{matrix}\right) & hi \\ my & name\end{matrix}\right] \left(\begin{matrix}is \\ slimshady\end{matrix}\right)\]

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

lmao